# Toward a Holistic Approach to Continual Model Merging

### v1

@@ Abstract
We present a holistic framework for continual model merging that intervenes at three critical stages: pre-merging, during merging, and post-merging—to address two fundamental challenges in continual learning. In particular, conventional approaches either maintain a growing list of per-domain task vectors, leading to scalability issues or rely solely on weight-space merging when old data is inaccessible, thereby losing crucial functional information. Our method overcomes these limitations by first fine-tuning the main model within its tangent space on domain-specific data; this linearization amplifies per-task weight disentanglement, effectively mitigating across-task interference. During merging, we leverage functional information from available optimizer states beyond mere parameter averages to avoid the need to revisit old data. Finally, a post-merging correction aligns the representation discrepancy between pre- and post-merged models, reducing bias and enhancing overall performance—all while operating under constant memory constraints without accessing historical data. Extensive experiments on standard class-incremental and domain-incremental benchmarks demonstrate that our approach not only achieves competitive performance but also provides a scalable and efficient solution to the catastrophic forgetting problem. Our code is available at https://github.com/VietHoang1512/CMM.
@@ 1 Introduction
Despite growing interest and wide application, existing merging techniques predominantly focus on the “during-merging” phase, where practitioners combine multiple models without access to their original training data [21, 22]. This limitation can lead to suboptimal performance, as naive merging approaches often overlook critical functional information embedded in the training process. Additionally, the number of stored models can grow linearly with the number of tasks, posing scalability challenges. To overcome these limitations, we propose a holistic framework that systematically addresses model merging across three key phases, pre-merging, during-merging, and post-merging. By leveraging task-specific data, our method exploits valuable functional information, leading to improved merging performance compared to conventional linear interpolation methods. Figure 1 highlights this improvement, illustrating accuracy and loss curves when interpolating between models trained on the first two tasks of the Cars [23] dataset, where our proposed method (pink) consistently outperforms linear interpolation (blue) across all settings.
@@ 5.2 Class-incremental experimental results
While EWC [73] improves over the naive baselines, its performance often lags behind more recent techniques. Note that our method consistently outperforms other constant-memory approaches and frequently matches or even surpasses the performance of methods with linear memory usage. For example, it achieves an average accuracy of 72.54%, outperforming RandMix, which reaches 70.4%. Also, ours achieves the best or second-best performance across all configurations, getting a $\approx 5\%$ margin over Zero-shot CLIP. On CIFAR100, our method achieves strong results across all incremental splits, confirming its robustness as the number of tasks grows. On ImageNet-R, a dataset designed for studying adversarial robustness and transfer learning, our method continues to maintain a leading position among non linear-memory approaches, showing resilience to distribution shifts. On the Cars dataset, our method improves performance by 2–3 percentage points over the second best method. These results demonstrate that our method not only adheres to strict memory constraints but also effectively retains knowledge across multiple tasks, making it a viable option for real-world applications that prioritize memory efficiency. In contrast, methods like MaxAbs show poor performance, even lagging behind LwF, despite requiring memory proportional to the number of tasks, $\Theta(N)$.
@@ 7 Conclusion
In this work, we tackle the challenges of applying model-merging techniques within the context of continual learning. Drawing on our findings, we propose a scalable and effective approach that provides a closed-form solution approximating the minimizer of the weighted task loss. Our extensive experimental results demonstrate the superiority of our method across various class-incremental and domain-incremental scenarios.

### v2

@@ Abstract
We present a holistic framework for Continual Model Merging (CMM) that intervenes at three critical stages: pre-merging, during merging, and post-merging—to address two fundamental challenges in continual learning. In particular, conventional approaches either maintain a growing list of per-domain task vectors, leading to scalability issues or rely solely on weight-space merging when old data is inaccessible, thereby losing crucial functional information. Our method overcomes these limitations by first fine-tuning the main model within its tangent space on domain-specific data; this linearization amplifies per-task weight disentanglement, effectively mitigating across-task interference. During merging, we leverage functional information from available optimizer states beyond mere parameter averages to avoid the need to revisit old data. Finally, a post-merging correction aligns the representation discrepancy between pre- and post-merged models, reducing bias and enhancing overall performance—all while operating under constant memory constraints without accessing historical data. Extensive experiments on standard class-incremental and domain-incremental benchmarks demonstrate that our approach not only achieves competitive performance but also provides a scalable and efficient solution to the catastrophic forgetting problem. Our code is available at https://github.com/VietHoang1512/CMM.
@@ 1 Introduction
Despite growing interest and wide application, existing merging techniques predominantly focus on the “during-merging” phase, where practitioners combine multiple models without access to their original training data [21, 22]. This limitation can lead to suboptimal performance, as naive merging approaches often overlook critical functional information embedded in the training process. Additionally, the number of stored models can grow linearly with the number of tasks, posing scalability challenges. To overcome these limitations, we propose CMM - a holistic framework that systematically addresses model merging across three key phases, pre-merging, during-merging, and post-merging. By leveraging task-specific data, our method exploits valuable functional information, leading to improved merging performance compared to conventional linear interpolation methods. Figure 1 highlights this improvement, illustrating accuracy and loss curves when interpolating between models trained on the first two tasks of the Cars [23] dataset, where our proposed method (pink) consistently outperforms linear interpolation (blue) across all settings.
@@ 5.2 Class-incremental experimental results
While EWC [73] improves over the naive baselines, its performance often lags behind more recent techniques. Note that our method consistently outperforms other constant-memory approaches and frequently matches or even surpasses the performance of methods with linear memory usage. For example, it achieves an average accuracy of 72.54%, outperforming RandMix, which reaches 70.4%. Also, CMM achieves the best or second-best performance across all configurations, getting a $\approx 5\%$ margin over Zero-shot CLIP. On CIFAR100, our method achieves strong results across all incremental splits, confirming its robustness as the number of tasks grows. On ImageNet-R, a dataset designed for studying adversarial robustness and transfer learning, our method continues to maintain a leading position among non linear-memory approaches, showing resilience to distribution shifts. On the Cars dataset, our method improves performance by 2–3 percentage points over the second best method. These results demonstrate that our method not only adheres to strict memory constraints but also effectively retains knowledge across multiple tasks, making it a viable option for real-world applications that prioritize memory efficiency. In contrast, methods like MaxAbs show poor performance, even lagging behind LwF, despite requiring memory proportional to the number of tasks, $\Theta(N)$.
@@ 7 Conclusion
In this work, we tackle the challenges of applying model-merging techniques within the context of continual learning. Drawing on our findings, we propose CMM - a scalable and effective approach that provides a closed-form solution approximating the minimizer of the weighted task loss. Our extensive experimental results demonstrate the superiority of our method across various class-incremental and domain-incremental scenarios.
@@
Acknowledgments
We thank anonymous reviewers for helpful feedback. QL acknowledges support of NSF DMS-2523382 and DOE Office of Science under Award #DE-SC0024721.